OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 4m = 4 + 4n Step 1: –4(m) = –4(8 + 2n) [Equation 1 is multiplied by –4.] 4m = 4 + 4n [Equation 2] Step 2: –4m = –32 – 8n [Equation 1 in Step 1 is simplified.] 4m = 4 + 4n [Equation 2] Step 3: –4m + 4m = –32 – 8n + 4n [Equations in Step 2 are added.] Step 4: 0 = –32 – 4n Step 5: n = –8 In which step did the student first make an error? Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

looks like the 4 got dropped in step 3 who wrote this please?

OpenStudy (anonymous):

this is the nonsense work i have to do. I don't know who wrote this by the way.

OpenStudy (anonymous):

is it an on line class?

OpenStudy (anonymous):

and if so, what system?

OpenStudy (anonymous):

ya and its brainhoney

Nnesha (nnesha):

Equation 1: m = 8 + 2n Equation 1: m = 8 + 2n try to find solution of these 2 equations don't look at the multiple choices yet if you know how to solve these type of questions then you willll be able to find mistakes

OpenStudy (anonymous):

@Lianna98 Bavel, brainhoney??? Module 8 practice test?